OpenStudy (anonymous):

The illumination I from a light varies inversely as the square of its distance d from the illuminated object. If I = 8.00 ft-candles when d = 5.00 ft, what is I when d = 4.00 ft

OpenStudy (anonymous):

I think it goes like this\[8=1\div d^2\] so \[8=200/5^2\] now substitute 4ft into the equation to give \[x = 200\div 4^2 =12.5\] any one have a better idea

OpenStudy (anonymous):

because this means the further away the less intensity there is